OpenStudy (helpmeplease14):

FAN AND MEDAL 10p-2(3p-6)=4(3p-6)-8p

OpenStudy (anonymous):

do we solve for p?

OpenStudy (helpmeplease14):

Yea

OpenStudy (anonymous):

PEMDAS wanna work it out? step by step?

OpenStudy (helpmeplease14):

I got 12=-24

OpenStudy (helpmeplease14):

HELP

OpenStudy (anonymous):

no we have to isolate (solve) for p, meaning p is on one side and the value on the other

OpenStudy (helpmeplease14):

Can you show me?

OpenStudy (anonymous):

10p-2(3p-6)=4(3p-6)-8p distribute -2 and 4 in the parenthesis 10p -6p+12=12p+18-8p combine like terms

OpenStudy (anonymous):

try it

OpenStudy (helpmeplease14):

Okay I did that

OpenStudy (helpmeplease14):

Combine -6p and -8p???

OpenStudy (anonymous):

10p -6p+12=12p+18-8p no not yet... 10p-6p=? 12p-8p=? you cannot do anything across the "=" yet

OpenStudy (helpmeplease14):

4p+12=4p-24

OpenStudy (anonymous):

wait wtf.. hold on... give me a minute? this don't make sense..

OpenStudy (helpmeplease14):

Okay

OpenStudy (anonymous):

what is the actual question, do you have a screenshot? because this honestly dont make sense right now

OpenStudy (helpmeplease14):

That's the question

OpenStudy (anonymous):

so it says solve for p? it doesn't ask you to do anything else?

OpenStudy (helpmeplease14):

No it just say solve each equation. Check your solution

OpenStudy (helpmeplease14):

And one of the problems is 10p-2(3p-6)=4(3p-6)-8p

OpenStudy (anonymous):

I dont see any solution to this, honestly..

OpenStudy (anonymous):

hold on, let me get a second opinion on this... @pooja195

OpenStudy (anonymous):

ive done it over a couple times... i see no solution.

pooja195 (pooja195):

\(\color{#0cbb34}{\text{Originally Posted by}}\) @atreyu6s3x6 ive done it over a couple times... i see no solution. \(\color{#0cbb34}{\text{End of Quote}}\) This is correct but next time guide the user as to why it is no solution

OpenStudy (anonymous):

okay thanks,, was just making sure i was correct.. @pooja195

OpenStudy (helpmeplease14):

So there's no solution??

OpenStudy (helpmeplease14):

Then 12=-24 is correct?

OpenStudy (anonymous):

but as you see your answer was correct and was done correctly.. @HelpMePlease14 there is no solution to this equation

OpenStudy (helpmeplease14):

Okay thank you

OpenStudy (anonymous):

no problem... haha thanks to the challenge xD